Opening stock 16000 Sales returns 250000 Purchases 150,000 Closing stock 25000 Staff costs 45000 Profit on exchange 1500 Electricity &water 25000 Profit sale shares 1000 Insurance 10000 Dividends from a wholly Rent and rates 5000 owned subsidiary 1000 Vehicle Expenses 7000 Interest received 500 NSSF 2000 Net loss 11500 Pensions and bonuses 5000 Bank charges and interest 1500 Subscriptions and donations 50 0 Legal Expenses 1000 Postage and telephone 1500 Bad debts 4000 Depreciation 5000 Other expenses 12000 TOTAL 290,50 0 290,500 The following further information is relevant a) Bad debts include an increase in provision of 1000 shillings b) Legal fees include 500 shilling relating to a directors private transaction c) Included in the vehicle expenses is a the cost of an old van valued at ksh 2000 d) Among the pension is 1500 paid to retired staff e) Interest received was from an overseas deposit account f) Included in the donations is ksh 500 to a children’s home calculate the adjusted profit for taxation


1
Expert's answer
2021-11-23T15:15:32-0500



Sales returns=250 000

Profit(highlighted in yellow)=1500+1000+5000+500+500=8500

expenses other positions

the adjusted (for the seller)=Profit-(Sales returns+expenses)=8500-(250 000+290 500+5500)=-537 500

the adjusted (for the buyers)=Profit+Sales returns-expenses=8500+250 000-290 500-5500)=-37 500

for the seller sales returns (-)

for the buyers sales returns (+)